Archie Co. purchased a framing machine for $45,000 on January 1, 2021. The machine is expected to have a four-year life, with a residual value of $5,000 at the end of four years.

Using the sum-of-the years'-digits method, depreciation for 2022 and book value at December 31, 2022, would be:

Multiple Choice

  • $13,500 and $13,500 respectively.

  • $13,500 and $8,500 respectively.

  • $12,000 and $17,000 respectively.

  • $12,000 and $12,000 respectively.
Answer #1

Depreciation for 2021 = (45,000-5,000)*4/10 = 16,000

Depreciation for 2022

= (45,000-5,000)*3/10 = 12,000

Book value = 45,000 - (16,000+12,000) = 17,000

Option C is the answer